# Solving the prisoner’s dilemma trap in Hamilton’s model of temporarily formed random groups

### v1

@@ 1 Introduction
Here, we show that the way we model reproduction can have a qualitative impact on the evolution of altruism in ways that the phenotypic approach cannot predict or explain. In particular, we show that strong altruism can evolve under biparental sexual reproduction in Hamilton’s model [2].
The rest of the paper is organized as follows. In Section 2 we describe Hamilton’s model of temporarily formed random groups, assuming $G$ groups of size $n$, using the terminology of the $n$-person prisoner’s dilemma. In Section 3 we develop an analytical treatment to solve Hamilton’s model in the deterministic limit $G\to\infty$ and show that in the case of asexual reproduction, cooperation is doomed to extinction in the strong altruism scenario, but not in the weak altruism scenario. In addition, we show that sexual reproduction, allowing self-mating, leads to the same conclusion as in the asexual case. We then consider biparental sexual reproduction and, challenging conventional wisdom, derive the conditions that guarantee the emergence of cooperation in the strong altruism scenario. In all cases, we follow the fate of the altruistic or selfish strategy, which for sexual reproduction would be analogous to following the allelic changes in the recombinant offspring chromosomes hosting the behavioral gene. In Section 4 we use simulations to prove the robustness of the conclusions of the deterministic analysis to the effect of demographic noise that arises when $G$ is finite. Finally, in Section 5 we summarize our results and also suggest a possible reassessment of cooperation in cases of viral co-infection, as cooperation may even be favored in situations where the prisoner’s dilemma applies. Returning to the phenotypic gambit, we cannot avoid ending this section by recalling a quote from Albert Einstein (possibly misattributed; see [21]): “Make things as simple as possible, but not simpler.”.
@@ 2 Model
In Hamilton’s model, within-group competition takes place during the reproduction stage. As mentioned before, we will consider three modes of reproduction. In the asexual reproduction mode, which is the one considered in Hamilton’s seminal paper, we select one parent within each group with probability equal to her relative fitness, and the single offspring inherits the strategy of the parent. For each group, exactly $m$ parents are selected in this way, and the resulting $m$ offspring are added to the offspring pool. Hamilton refers to the offspring pool as the migrant pool, which is formed by young individuals that reach maturity [2]. In sexual reproduction with selfing allowed, we randomly select two not necessarily different parents, and the single offspring inherits the strategy of either parent with equal probability. This selection process is repeated $m$ times resulting in the addition of $m$ offspring to the offspring pool. This mode is clearly equivalent to asexual reproduction, as we will show in the following section. Finally, in the biparental sexual reproduction mode, the two randomly chosen parents must be different and, as before, the single offspring inherits the strategy of either parent with equal probability. In this mode, it is guaranteed that in a group with $i=n-1$ cooperators, one cooperator will be chosen as a parent. Note that the groups contribute different numbers of offspring to the offspring pool, not only because $m$ is a Poisson distributed random variable, but also because the parameter of the Poisson, i.e. $\pi(i)$, depends on the number of cooperators within the group, which varies from group to group. Finally, each group of the next generation is formed by sampling $n$ individuals with replacement from the offspring pool, and the whole process is repeated until one of the strategies disappears and evolution stops.
@@ 3.2 Sexual reproduction with self-mating
In this case, we randomly select two parents according to their relative fitness, and the single offspring inherits the strategy of either parent with equal probability. Since there are no restrictions on the choice of parents, the probability that an offspring is a cooperator is
@@ 3.3 Biparental sexual reproduction
Figure 1: Phase diagram for $n=4$ showing the stability regions of the different fixed points. In region C, only the all-cooperators fixed point $q^{*}=1$ is stable, while in region D, only the all-defectors fixed point $q^{*}=0$ is stable. In region C+D both fixed points are stable. The regions C and C+D are separated by the curve $r_{s}=r_{s}(c)$ and the regions D and C+D are separated by the curve $r_{a}=r_{a}(c)$.
The results of the stability analysis are conveniently summarized in a phase diagram in the plane $(c,r)$ as shown in Fig.  1 for $n=4$. Only the fixed point $q^{*}=1$ is stable in region C and only the fixed point $q^{*}=0$ is stable in region D. In region C+D we have the bistability situation, where the outcome of the dynamics depends on the initial frequency of cooperators $q_{0}$. In particular, the unstable coexistence fixed point $10<q^{*}<$ 1 determines the domains of attraction of the two stable fixed points: the dynamics leads to the all-cooperators regime when $q_{0}>q^{*}$ and to the all-defectors regime when $q_{0}<q^{*}$. An important feature of this phase diagram, which also holds for general group sizes $n>2$, is that there is no stable coexistence solution. In fact, a stable coexistence fixed point would occur if $r_{s}<r_{a}$ in the limit $c\to 0$, so there would be a region where neither of the extreme fixed points is stable, but we have analytically ruled out this possibility by showing that $r_{a}\to r_{s}$ in this limit.
@@
For large $n$ we have $r_{a}\approx 2n^{2}$, and for the group size $n=4$ used in drawing the phase diagram in Fig. 1 we have $r_{a}\approx 10.29$.
Figure 2: Unstable coexistence fixed point $q^{*}$ as a function of the amplification factor $r$ for $c=1$ and (from left to right) $n=3,4,5,6$. If $q_{0}>q^{*}$ the dynamics is driven to the all-cooperators fixed point and to the all-defectors fixed point if $q_{0}<q^{*}$. The value of $r$ at which $q^{*}=1$ signals the instability of the all-cooperators fixed point and is given by the solution of Eq. (30). The all-defectors fixed point is stable for these parameter settings.
@@ 5 Discussion
There is a consensus that strong altruism, where a cooperator does not benefit from her costly altruistic behavior, cannot evolve in randomly formed, transient groups unless there is some sort of positive assortment among cooperators [2, 3, 4, 24, 32, 33] or punishment for defectors [10, 15, 16]. Note that reputation and reciprocity foster cooperation through the creation of positive assortment among players [34]. Even with these fixes, one difficulty remains: the so-called ‘origination problem’, i.e. how cooperation can take off from an initial low frequency of cooperators. However, it has been argued that this problem is an artifact of assuming, as ours and most models do, that strategies exist as discrete categories rather than as quantitative traits, as is most likely the case [29]. Alas, the quantitative strategy trait fix has only been tested in the case of weak altruism, where a cooperator also benefits from her altruistic behavior and the benefit exceeds the cost of cooperation [29]. In this case, even assuming discrete traits, there are conditions under which defectors can be invaded by rare cooperators, as we showed in Section 3.1. For the others-only variant of the $n$-prisoner’s dilemma, and for scenarios of strong altruism in general, the problem of origination remains.
@@
The main purpose of this paper was to highlight the importance that the mode of reproduction can have on the fate of strong altruism in populations that are frequently divided into small temporary groups, in which individuals interact locally and fitness-related processes such as mating take place. The selection dynamics resulting from the different forms of reproduction considered here were solved analytically in the limit of infinitely many groups, assuming, as usual, that individuals colonize the groups according to a binomial distribution, so that cooperators do not cluster into groups, and that local interactions follow the others-only variant of the $n$-prisoner’s dilemma. Under biparental sexual reproduction, the results are quite striking, challenging conventional wisdom by showing that strong altruism can indeed evolve in the absence of assortment and punishment in temporary groups if the amplification factor $r$ is sufficiently large. The basic reason is that biparental sexual reproduction may enable a resident population of cooperators to resist invasion by rare defectors, since the fitness advantage of a single defector in a group with $n-1$ cooperators is smothered by the requirement that the recombinant offspring chromosome has a one-half chance of being a cooperator in any mating involving the defector. However, and most surprisingly, biparental sexual reproduction can also enable the invasion of a resident population of defectors by rare cooperators if the cost of cooperation $c$ is not large, thus solving the problem of origination. These results contrast strikingly with asexual reproduction and sexual reproduction with self-mating modes, for which cooperation is always doomed and the amplification factor $r$ can only slow the rate at which cooperators disappear from the population [2].
@@
These findings are summarized in Fig. 1, which shows a region, labeled C, where cooperation is the only evolutionary outcome. For large group size $n$, this region exists for $c<1/2n$ and $r>2n^{2}$. The only modification we have made to Hamilton’s original model [2] is to change asexual reproduction to biparental sexual reproduction, which is how most sexual reproduction occurs [20]. This result provides a neat solution to both the problem of origination and the conditions under which strong altruism can evolve, thus avoiding the ‘inverse genetic fallacy’ [35], i.e. the inappropriate attribution of mechanisms that might support cooperation (e.g., positive assortment and punishment) to the explanation of its origin. For large cooperation costs (i.e., $c\approx 1$) and large $n$, we find a wide region in the parameter space, labeled C+D in Fig. 1, bounded at the bottom by $r=2n^{2}$, where cooperation can evolve provided that $q_{0}>4n/r$, where $q_{0}$ is the initial frequency of cooperators. These estimates for large $n$ are worst-case scenarios for cooperation, since increasing group size $n$ favors defection. In addition, we have shown that the results of our analytical study for infinitely many groups are robust to the demographic noise that arises when the number of groups is finite.
@@
Finally, it is an open question whether our conclusions apply to diploid sexual organisms under the more realistic assumption that altruistic and selfish strategies are quantitative traits rather than discrete behaviors. As Maynard Smith wrote [43]: “Most populations of interest have sexual diploid inheritance. In many cases this does not matter.” Apart from the obvious case where a pure phenotypic strategy is produced by a heterozygous genotype and cannot be fixed in a random mating population because of Mendelian inheritance, we believe there is no simple answer to the question of whether it does or does not matter.

### v2

@@ 1 Introduction
Here, we show that the way we model reproduction can have a qualitative impact on the evolution of altruism in ways that the phenotypic approach cannot predict or explain. In particular, we show that strong altruism can evolve under biparental sexual reproduction in Hamilton’s model [2]. Strictly speaking, and considering eukaryotes, we analyzed the fate of altruism in a monoecious population, where individuals have both male and female reproductive organs and can produce gametes of both sexes, but cannot self-fertilize. Earthworms, for example, are monoecious. This nondifferentiation between the sexes simplifies the model and allows a full analytical solution in the infinite population limit. In an appendix, we use simulations to extend the analyses to a dioecious population (with distinct unisexual individuals) and show that sex differentiation actually promotes altruistic traits.
The rest of the paper is organized as follows. In Section 2 we describe Hamilton’s model of temporarily formed random groups, assuming $G$ groups of size $n$, using the terminology of the $n$-person prisoner’s dilemma. In Section 3 we develop an analytical treatment to solve Hamilton’s model in the deterministic limit $G\to\infty$ and show that in the case of asexual reproduction, cooperation is doomed to extinction in the strong altruism scenario, but not in the weak altruism scenario. In addition, we show that sexual reproduction, allowing self-mating, leads to the same conclusion as in the asexual case. We then consider biparental sexual reproduction and, challenging conventional wisdom, derive the conditions that guarantee the emergence of cooperation in the strong altruism scenario. In all cases, we follow the fate of the altruistic or selfish strategy, which for sexual reproduction would be analogous to following the allelic changes in the recombinant offspring chromosomes hosting the behavioral gene. In Section 4 we use simulations to prove the robustness of the conclusions of the deterministic analysis to the effect of demographic noise that arises when $G$ is finite. Finally, in Section 5 we summarize our results and also suggest a possible reassessment of cooperation in cases of viral co-infection, as cooperation may even be favored in situations where the prisoner’s dilemma applies. In A we present the simulations for a dioecious population with distinct unisexual individuals. Returning to the phenotypic gambit, we cannot avoid ending this section by recalling a quote from Albert Einstein (possibly misattributed; see [21]): “Make things as simple as possible, but not simpler.”.
@@ 2 Model
In Hamilton’s model, within-group competition takes place during the reproduction stage. As mentioned before, we will consider three modes of reproduction. In the asexual reproduction mode, which is the one considered in Hamilton’s seminal paper, we select one parent within each group with probability equal to her relative fitness, and the single offspring inherits the strategy of the parent. For each group, exactly $m$ parents are selected in this way, and the resulting $m$ offspring are added to the offspring pool. Hamilton refers to the offspring pool as the migrant pool, which is formed by young individuals that reach maturity [2]. In sexual reproduction with selfing allowed, we randomly select two not necessarily different parents with probability given by their relative fitness, and the single offspring inherits the strategy of either parent with equal probability. This selection process is repeated $m$ times resulting in the addition of $m$ offspring to the offspring pool. This mode is clearly equivalent to asexual reproduction, as we will show in the following section. Finally, in the biparental sexual reproduction mode, the two parents, randomly chosen with probability given by their relative fitness, must be different, and the single offspring, as before, inherits the strategy of either parent with equal probability. In this mode, it is guaranteed that in a group with $i=n-1$ cooperators, one cooperator will be chosen as a parent. Note that the groups contribute different numbers of offspring to the offspring pool, not only because $m$ is a Poisson distributed random variable, but also because the parameter of the Poisson, i.e. $\pi(i)$, depends on the number of cooperators within the group, which varies from group to group. Finally, each group of the next generation is formed by sampling $n$ individuals with replacement from the offspring pool, and the whole process is repeated until one of the strategies disappears and evolution stops.
@@ 3.2 Sexual reproduction with self-mating
In this case, we randomly select two parents with replacement according to their relative fitness, and the single offspring inherits the strategy of either parent with equal probability. Since there are no restrictions on the choice of parents, the probability that an offspring is a cooperator is
@@ 3.3 Biparental sexual reproduction
Figure 1: Phase diagram for $n=4$ (left panel) and $n=8$ (right panel) showing the stability regions of the different fixed points. In region C, only the all-cooperators fixed point $q^{*}=1$ is stable, while in region D, only the all-defectors fixed point $q^{*}=0$ is stable. In region C+D both fixed points are stable. The regions C and C+D are separated by the curve $r_{s}=r_{s}(c)$ and the regions D and C+D are separated by the curve $r_{a}=r_{a}(c)$.
The results of the stability analysis are conveniently summarized in a phase diagram in the plane $(c,r)$ as shown in Fig.  1 for $n=4$ and $n=8$. Only the fixed point $q^{*}=1$ is stable in region C and only the fixed point $q^{*}=0$ is stable in region D. In region C+D we have the bistability situation, where the outcome of the dynamics depends on the initial frequency of cooperators $q_{0}$. In particular, the unstable coexistence fixed point $10<q^{*}<$ 1 determines the domains of attraction of the two stable fixed points: the dynamics leads to the all-cooperators regime when $q_{0}>q^{*}$ and to the all-defectors regime when $q_{0}<q^{*}$. An important feature of these phase diagrams, which also holds for general group sizes $n>2$, is that there is no stable coexistence solution. In fact, a stable coexistence fixed point would occur if $r_{s}<r_{a}$ in the limit $c\to 0$, so there would be a region where neither of the extreme fixed points is stable, but we have analytically ruled out this possibility by showing that $r_{a}\to r_{s}$ in this limit.
@@
For large $n$ we have $r_{a}\approx 2n^{2}$, and for the group sizes used in drawing the phase diagrams in Fig. 1 we have $r_{a}\approx 10.29$ for $n=4$ and $r_{a}\approx 78.09$ for $n=8$.
Figure 2: Unstable coexistence fixed point $q^{*}$ as a function of the amplification factor $r$ for $c=1$ and (from left to right) $n=3,4,5,6$. If $q_{0}>q^{*}$ the dynamics is driven to the all-cooperators fixed point and to the all-defectors fixed point if $q_{0}<q^{*}$, as indicated by the arrows. The value of $r$ at which $q^{*}=1$ signals the instability of the all-cooperators fixed point and is given by the solution of Eq. (30). The all-defectors fixed point is stable for these parameter settings.
@@ 5 Discussion
There is a consensus that strong altruism, where a cooperator does not benefit from her costly altruistic behavior, cannot evolve in randomly formed, transient groups unless there is some sort of positive assortment among cooperators [2, 3, 4, 24, 32, 33] or punishment for defectors [10, 15, 16]. Note that reputation and reciprocity foster cooperation through the creation of positive assortment among players [34]. Even with these fixes, one difficulty remains: the so-called ‘origination problem’, i.e. how cooperation can take off from an initial low frequency of cooperators.
A word on the terminology of weak and strong altruism is in order. Whether we are dealing with weak or strong altruism is determined by the fitness of the altruist or cooperator, Eq. (1): if the benefit she receives from her own contribution is greater than or equal to the cost of the altruistic act, then we have weak altruism, otherwise we have strong altruism. This terminology considers only first-order effects, the direct effect of the altruist on herself. There are also second-order or indirect effects, where the altruist helps others who in turn help the altruist. The various prescriptions for the emergence of cooperation in the strong altruism scenario (e.g., positive assortment, punishment, and reputation) involve playing with these second-order effects. Of course, if such a prescription works, it is because the benefits of the second-order effects exceed the costs of the altruistic acts. Thus, including these indirect effects in the benefits accruing to the altruist leads to the conclusion that, by definition, cooperation cannot evolve in the strong altruism scenario.
@@
The main purpose of this paper was to highlight the importance that the mode of reproduction can have on the fate of strong altruism in populations that are frequently divided into small temporary groups, in which individuals interact locally and fitness-related processes such as mating take place. The selection dynamics resulting from the different forms of reproduction considered here were solved analytically in the limit of infinitely many groups, assuming, as usual, that individuals colonize the groups according to a binomial distribution, so that cooperators do not cluster into groups, and that local interactions follow the others-only variant of the $n$-prisoner’s dilemma. Under biparental sexual reproduction, the results are quite striking, challenging conventional wisdom by showing that strong altruism can indeed evolve in the absence of assortment and punishment in temporary groups if the amplification factor $r$ is sufficiently large. The basic reason is that biparental sexual reproduction may enable a resident population of cooperators to resist invasion by rare defectors, since the fitness advantage of a single defector in a group with $n-1$ cooperators is smothered by the requirement that the recombinant offspring chromosome has a one-half chance of being a cooperator in any mating involving the defector. However, and most surprisingly, biparental sexual reproduction can also enable the invasion of a resident population of defectors by rare cooperators if the cost of cooperation $c$ is not large, thus solving the problem of origination. These results contrast strikingly with asexual reproduction and sexual reproduction with self-mating modes, for which cooperation is always doomed and the amplification factor $r$ can only slow the rate at which cooperators disappear from the population [2]. Although our conclusions apply strictly to populations of monoecious individuals, where each individual can produce both male and female gametes but they cannot self-fertilize, they also apply to dioecious populations, where individuals can produce either male or female gametes (see A). In particular, we find that dioecy promotes cooperation.
@@
Our main findings are summarized in Fig. 1, which shows a region, labeled C, where cooperation is the only evolutionary outcome. For large group size $n$, this region exists for $c<1/2n$ and $r>2n^{2}$. The only modification we have made to Hamilton’s original model [2] is to change asexual reproduction to biparental sexual reproduction, which is how most sexual reproduction occurs [20]. This result provides a neat solution to both the problem of origination and the conditions under which strong altruism can evolve, thus avoiding the ‘inverse genetic fallacy’ [35], i.e. the inappropriate attribution of mechanisms that might support cooperation (e.g., positive assortment and punishment) to the explanation of its origin. For large cooperation costs (i.e., $c\approx 1$) and large $n$, we find a wide region in the parameter space, labeled C+D in Fig. 1, bounded at the bottom by $r=2n^{2}$, where cooperation can evolve provided that $q_{0}>4n/r$, where $q_{0}$ is the initial frequency of cooperators. These estimates for large $n$ are worst-case scenarios for cooperation, since increasing group size $n$ favors defection. In addition, we have shown that the results of our analytical study for infinitely many groups are robust to the demographic noise that arises when the number of groups is finite.
@@
A noticeable feature of our study is the high values of the amplification factor $r$, which measures the benefit of the cooperators to the group, necessary for the emergence of cooperation in the scenario of biparental sexual reproduction. This benefit is the mean number of offspring produced by the group, which grows linearly with $r$ according to Eqs. (3) and (7). The population biology of early replicators, in particular the enzyme production problem [43, 44], provides a scenario where $r$ is expected to be very large. Here the cooperators are enzyme producers and the defectors benefit from the enzyme but do not encode it (so the defectors are shorter than the cooperators and therefore replicate faster). The fitness used to model this problem is exactly the weak altruism version of the n-person prisoner’s dilemma, and since enzyme-catalyzed reactions are often many orders of magnitude faster than their non-catalyzed counterparts, it follows that $r$ must be very large [43]. In fact, we note that the interaction between enzyme and replicator is a hypercyclic coupled biosynthetic process, and the classical experiments on phage-infected bacteria show that during the early stage of viral infection there is a hyperbolic increase in the viral RNA synthesis rate, which is a clear sign of hypercyclic coupling [45, 46]. This suggests that the role of hypercyclic coupling in early molecular evolution is akin to its role in viral evolution today, which, as discussed above, provides a biological context for our model.
Finally, it is an open question whether our conclusions apply to diploid sexual organisms under the more realistic assumption that altruistic and selfish strategies are quantitative traits rather than discrete behaviors. Indeed, it has been argued that if strategies exist as quantitative traits, rather than as discrete traits as ours and most models assume, then there are conditions under which defectors can be invaded by rare cooperators, at least in a weak altruism scenario [29]. As Maynard Smith wrote [47]: “Most populations of interest have sexual diploid inheritance. In many cases this does not matter.” Apart from the obvious case where a pure phenotypic strategy is produced by a heterozygous genotype and cannot be fixed in a random mating population because of Mendelian inheritance, we believe there is no simple answer to the question of whether it does or does not matter.
@@ Acknowledgments
Appendix A
In the main text we considered a population composed of monoecious individuals, i.e. individuals that have both male and female reproductive organs, and can produce gametes of both sexes. Here we consider a finite population of dioecious individuals, so that only males and females can mate. We also assume equal sex ratio and no correlation between sex and strategies. The assumption of equal sex ratios applies to the population as a whole, not to the sex ratio in a particular group, which may vary according to binomial sampling. The model is considerably more complicated than the monoecious case discussed in the main text, so we limit our analysis to simulations of finite populations, which is sufficient to show that our main conclusions also hold in dioecious populations. Surprisingly, we find that cooperation is more likely to evolve in a population of dioecious individuals than in a population of monoecious individuals.
First, all $G\times n$ individuals are randomly assigned a strategy (cooperator or defector) and a sex (male or female). Then, $G$ groups of size $n$ are formed by randomly selecting individuals without replacement. Fitness do not depend on the sex of the individuals so the fitness of a cooperator in a group with $i-1$ cooperators $f_{c}(i)$ is given by Eq. (1), the fitness of a defector in a group with $i$ cooperators $f_{d}(i)$ is given by Eq. (2), and the mean fitness of a group $\pi(i)$ with $i$ cooperators is given by Eq. (3). The number of offspring $m$ produced by a group with $i$ cooperators is a Poisson random variable with mean $\pi(i)$ (see Eq. 5) if the group contains individuals of both sexes, otherwise $m=0$. This is the first major difference from the monoecious case, discussed in the main text. The second major difference, of course, is the choice of parents, who must now be of different sexes.
The strategy of an offspring is determined by the strategies of their parents. First, we choose a father from among the males in the group with a probability given by his relative fitness. A male’s relative fitness is given by the ratio of his fitness to the total fitness of the males in the group, so the choice of a father involves only competition between males. Then we choose a mother from among the females in the group with probability proportional to her fitness, following the same procedure described above: the choice of a mother involves a competition between females only. The offspring will inherit either the strategy of the father or the strategy of the mother with equal probability. To avoid correlations between sex and strategy, the sex of the offspring is randomly determined with equal probability.
Figure 1: Probability of fixation of the cooperators as a function of the amplification factor $r$ for (from left to right) $n=3,4,5,6,7$ (left panel). The right panel shows the rough estimate of the threshold $r_{c}$ obtained with the condition $\Pi=0.5$. The line is the fit $r_{c}=0.91-3.63n+7.20n^{2}$. The other parameters are $G=400$ and $c=1$.
The offspring pool is formed by the offspring produced by each of the $G$ groups, and the new groups are formed by sampling with replacement $G\times n$ individuals from this pool, as done for the population of monoecious individuals. The procedure is then repeated until one of the strategies becomes fixed. The left panel of Fig. 1 shows the probability of fixation of the cooperators for different group sizes $n$, which should be compared with the left panel of Fig. 6. The remarkable result is that the transition to the all-cooperators regime occurs for a smaller value of the amplification factor $r$ compared to the results for the monoecious individuals. A rough estimate of the threshold $r_{c}$, obtained by finding the value of $r$ at which $\Pi=0.5$, is shown in the right panel of Fig. 1. The results show that $r_{c}$ increases with $n^{2}$, similar to the case of biparental sexual reproduction discussed in the main text. We have also found that the all-cooperators regime is the only absorbing state for $n=2$, as expected. In fact, as argued before, the only advantage of a defector is to share the group with a cooperator, but this leverage is lost because the offspring have a 50% chance of inheriting the cooperative trait.
Why is the threshold $r_{c}$ lower for dioecious individuals? To gain some insight into this, we recall that increasing group size hinders cooperation. Since we restrict intragroup competition only between males and between females, the effective group size perceived by each group member is reduced, which explains our results.